# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ 3.2 Special Cases of the Bregman Divergence
  • Squared distance (squared loss): $g^{\text{SQ}}(\alpha)\coloneqq(\alpha-C)^{2}/2$ for some constant $C\in{\mathbb{R}}$.
  • Unnormalized KL (UKL) divergence: $g^{\text{UKL}}(\alpha)\coloneqq\big(|\alpha|-C\big)\log\big(|\alpha|-C\big)-|\alpha|$ for some constant $C<\inf_{x}\alpha(x)$.
  • Binary KL (BKL) divergence: $g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-C)\log\big(|\alpha|-C\big)-(|\alpha|+C)\log(|\alpha|+C)$ for some constant $C<\inf_{x}\alpha(x)$.
  • Basu’s power (BP) divergence (BP-Riesz): $g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-C\big)^{1+\gamma}-\big(|\alpha|-C\big)}{\gamma}-\big(|\alpha|-C\big)$ for some $\gamma\in(0,\infty)$ and $C<\inf_{x}\alpha(x)$.
@@ 3.3 SQ-Riesz Regression
$\partial g^{\text{SQ}}(\alpha)$ $=2\big(\alpha-C\big).$
@@
$\text{BD}_{g}\big(\alpha\big)\coloneqq\mathbb{E}\Big[\alpha(X)^{2}-m\big(W,2\big(\alpha(\cdot)-C\big)\big)\Big].$
@@
$\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha(X_{i})^{2}-2m\big(W_{i},\big(\alpha(\cdot)\big)\big)\Big).$
Here, for simplicity, we drop constant terms that are irrelevant for the optimization and use the linearity of $m$ for $2(\alpha(\cdot)-C)$ [^1] The original empirical Bregman divergence objective using $g(\alpha)=(\alpha-C)^{2}$ in (1) is given as $\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)$ $={\mathbb{E}}\left[-(\alpha(X)-C)^{2}+2(\alpha(X)-C)\alpha(X)-m\big(W,2(\alpha(\cdot)-C)\big)\right]$ $={\mathbb{E}}\left[\alpha(X)^{2}-C^{2}-m\big(W,2\big(\alpha(\cdot)-C\big)\big)\right].$ . This estimation method corresponds to Riesz regression in debiased machine learning (Chernozhukov et al., 2021) and least-squares importance fitting (LSIF) in density ratio estimation (Kanamori et al., 2009). Moreover, if we define $\mathcal{A}$} appropriately, we can recover nearest neighbor matching, as pointed out in Kato (2025a), which extends the argument in Lin et al. (2023).
@@ 3.4 UKL-Riesz Regression
Next, we consider a KL-divergence-motivated convex function. Let $C<\inf_{x}\alpha(x)$ be a constant. We define
@@
Under this choice of $g$, the Bregman divergence objective is given as follows [^2] This Bregman divergence objective is derived as follows: $\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)$ $={\mathbb{E}}\Big[-\left[|\alpha(X)|-C\right])\log\left(|\alpha(X)|-C\right)-|\alpha|+\operatorname{sign}\left(\alpha(X)\right)\alpha(X)\log\left(|\alpha(X)|-C\right)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\Big(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\Big].$ :
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)\coloneqq\mathbb{E}\Big[C\log\left(|\alpha(X_{i})|-C\right)-|\alpha(X_{i})|-m\Big(W_{i},\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\Big].$
@@
$\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)$ $=\frac{1}{n}\sum^{n}_{i=1}\left(C\log\left(|\alpha(X_{i})|-C\right)-|\alpha(X_{i})|-m\Big(W_{i},\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\right).$
@@ 3.5 BKL-Riesz Regression
We introduce the BKL divergence and BKL-Riesz, which are motivated by the KL divergence and logistic regression. Let $C<\inf_{x}\alpha(x)$ be a constant. We define
@@ 3.6 BP-Riesz Regression
Basu’s power (BP) divergence bridges the squared loss and KL divergence (Basu et al., 1998). Let $C<\inf_{x}\alpha(x)$ be a constant. Based on the BP divergence, we introduce the following function:
$g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-C\big)^{1+\gamma}-\big(|\alpha|-C\big)}{\gamma}-|\alpha|.$
@@
$\partial g^{\text{BP}}(\alpha)=\left(1+\frac{1}{\gamma}\right)\operatorname{sign}(\alpha)\Big(\big(|\alpha|-C\big)^{\gamma}-1\Big).$
@@
Under this choice of $g$, the Bregman divergence objective is given as follows [^4] This Bregman divergence objective is derived from $\text{BD}_{g^{\text{BP}}}\big(\alpha\big)$ $\coloneqq{\mathbb{E}}\Bigg[-\frac{\big(|\alpha(X)|-C\big)^{1+\gamma}-\big(|\alpha(X)|-C\big)}{\gamma}+|\alpha(X)|+(1+1/\gamma)|\alpha(X)|\Big(\big(|\alpha(X)|-C\big)^{\gamma}-1\Big)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,(1+1/\gamma)\operatorname{sign}(\alpha)\Big(\big(|\alpha(\cdot)|-C\big)^{\gamma}-1\Big)\right)\Bigg].$ :
$\text{BD}_{g^{\text{BP}}}\big(\alpha\big)\coloneqq{\mathbb{E}}\Bigg[\frac{C\big(\big(|\alpha(X)|-C\big)^{\gamma}-1\big)}{\gamma}+C\big(|\alpha(X)|-C\big)^{\gamma}+\big(|\alpha(X)|-C\big)^{1+\gamma}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,\left(1+\frac{1}{\gamma}\right)\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-C\big)^{\gamma}-1\Big)\right)\Bigg].$
@@
$\widehat{\text{BD}}_{g^{\text{BP}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Bigg(\frac{C\big(\big(|\alpha(X_{i})|-C\big)^{\gamma}-1\big)}{\gamma}+C\big(|\alpha(X_{i})|-C\big)^{\gamma}+\big(|\alpha(X_{i})|-C\big)^{1+\gamma}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W_{i},\left(1+\frac{1}{\gamma}\right)\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-C\big)^{\gamma}-1\Big)\right)\Bigg).$
Basu’s power divergence bridges the squared loss and the (U)KL divergence. When $\gamma=1$, BP-Riesz regression reduces to SQ-Riesz regression, while when $\gamma\to 0$, BP-Riesz regression reduces to UKL-Riesz regression. This follows because
$\lim_{\gamma\to 0}\frac{\big(|\alpha|-C\big)^{\gamma}-1}{\gamma}=\log\big(|\alpha|-C\big).$
@@ 3.7 PU-Riesz Regression
We introduce PU learning loss and PU-Riesz, which are motivated by PU learning. Let $C<\inf_{x}\alpha(x)$ be some constant. We define $g^{\text{PU}}$ as
@@
PU learning is a classical problem. For example, Lancaster & Imbens (1996) studies this problem under a stratified sampling scheme (Wooldridge, 2001). du Plessis et al. (2015) refinds this formulation and calls it unbiased PU learning. Kato et al. (2019) points out the relationship between PU learning and density ratio estimation, and Kato & Teshima (2021) shows that PU learning is a special case of density ratio model fitting under a Bregman divergence. Our results further generalize these results. Note that PU learning in these settings and our setting is called case-control PU learning. There is also another formulation called censoring PU learning (Elkan & Noto, 2008). Kato et al. (2025) considers ATE estimation in a PU learning setup and applies our method in their study.
@@ 4 Automatic Covariate Balancing
Under specific choices of the Riesz representer model and the Bregman divergence, we can guarantee an automatic covariate balancing property. The key tool is the duality between Bregman divergence minimization and covariate balancing methods.
@@ 4.2 Examples
We can improve estimation accuracy by incorporating additional modeling assumptions. For example, in ATE estimation, we can approximate the propensity score by a logistic model,
$e_{{\bm{\beta}}}(X)\coloneqq\frac{1}{1+\exp\Big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)}.$
Plugging this propensity score model into the Riesz representer for ATE, we can approximate the Riesz representer by
$\alpha^{\text{ATE}}_{{\bm{\beta}}}(X)\coloneqq\frac{D}{e_{{\bm{\beta}}}(X)}+\frac{1-D}{1-e_{{\bm{\beta}}}(X)}.$
@@
$\zeta^{-1}\left(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\right)$ $=\frac{D}{e_{{\bm{\beta}}}(X)}+\frac{1-D}{1-e_{{\bm{\beta}}}(X)}$
$=D\Big(1+\exp\Big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)\Big)+(1-D)\Big(1+\exp\Big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)\Big).$
@@ 4.3 Automatic Covariate Balancing
This section presents a covariate balancing property for the basis function ${\bm{\phi}}(X)$ when the Riesz representer model has the form $\alpha(X)=\zeta^{-1}\big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)$. Related properties have been reported in Ben-Michael et al. (2021). More specifically, for Riesz regression, Bruns-Smith & Feller (2022) and Bruns-Smith et al. (2025) discuss the duality between Riesz regression and stable balancing weights (Zubizarreta, 2015). Zhao (2019) also reports the duality between tailored loss minimization and entropy balancing weights. Tarr & Imai (2025) investigates such a duality in covariate balancing method using the support vector machine (SVM). In density ratio estimation, such a duality is also well known, as discussed in Sugiyama et al. (2007) and Sugiyama et al. (2011b). In the density ratio estimation literature, the resulting balancing property is typically interpreted as moment matching. This study generalizes these results and provides new findings on the balancing property.
@@
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{k}_{j=1}\beta_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right);$
that is, $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)$ is linear in ${\bm{\phi}}(X_{i})^{\top}{\bm{\beta}}$. In this case, a Riesz representer model $\widehat{\alpha}=\alpha_{\widehat{{\bm{\beta}}}}$ trained by empirical risk minimization satisfies
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\widetilde{g}\big(X_{i},\phi_{j}(X_{i})\big)-m(W_{i},\phi_{j})\Big)\right|\leq\lambda.$
@@
This result corresponds to covariate balancing in ATE estimation and moment matching in density ratio estimation. In ATE estimation, if we use a linear link and SQ-Riesz regression, we have
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}_{i}\phi_{j}(D_{i},Z_{i})-\frac{1}{n}\sum^{n}_{i=1}\Big(\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\Big)\right|<\lambda\quad(\text{for}\ j=1,2,\dots,p).$
@@
$\left|\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\alpha_{i}\phi_{j}(1,Z_{i})-\mathbbm{1}[D_{i}=0]\alpha_{i}\phi_{j}(0,Z_{i})\Big)\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
@@
$\left|\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-\sum^{n}_{i=1}\phi_{j}(X_{i})\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
@@
$\left|\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-1\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$

### v2

@@ 3.2 Special Cases of the Bregman Divergence
  • Squared distance (squared loss): $g^{\text{SQ}}(\alpha)\coloneqq(\alpha-C)^{2}$ for some constant $C\in{\mathbb{R}}$.
  • Unnormalized KL (UKL) divergence: $g^{\text{UKL}}(\alpha)\coloneqq\big(|\alpha|-C\big)\log\big(|\alpha|-C\big)-|\alpha|$ for some constant $C<\inf_{x}|\alpha(x)|$.
  • Binary KL (BKL) divergence: $g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-C)\log\big(|\alpha|-C\big)-(|\alpha|+C)\log(|\alpha|+C)$ for some constant $C<\inf_{x}|\alpha(x)|$.
  • Basu’s power (BP) divergence (BP-Riesz): $g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-C\big)^{1+\delta}-\big(|\alpha|-C\big)}{\delta}-\big(|\alpha|-C\big)$ for some $\delta\in(0,\infty)$ and $C<\inf_{x}|\alpha(x)|$.
@@ 3.3 SQ-Riesz Regression
$\partial g^{\text{SQ}}(\alpha)$ $=2(\alpha-C).$
@@
$\text{BD}_{g}\big(\alpha\big)\coloneqq\mathbb{E}\Big[\alpha(X)^{2}-2m\big(W,\big(\alpha(\cdot)-C\big)\big)\Big].$
@@
$\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\alpha(X_{i})^{2}-2m\big(W_{i},\big(\alpha(\cdot)\big)\big)\right).$
Here, for simplicity, we drop constant terms that are irrelevant for the optimization and use the linearity of $m$ for $2(\alpha(\cdot)-C)$ [^1] The original empirical Bregman divergence objective using $g(\alpha)=(\alpha-C)^{2}$ in (1) is given as $\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)$ $={\mathbb{E}}\left[-(\alpha(X)-C)^{2}+2(\alpha(X)-C)\alpha(X)-2m\big(W,(\alpha(\cdot)-C)\big)\right]$ $={\mathbb{E}}\left[\alpha(X)^{2}-C^{2}-2m\big(W,\big(\alpha(\cdot)-C\big)\big)\right].$ . This estimation method corresponds to Riesz regression in debiased machine learning (Chernozhukov et al., 2021) and least-squares importance fitting (LSIF) in density ratio estimation (Kanamori et al., 2009). Moreover, if we define $\mathcal{A}$} appropriately, we can recover nearest neighbor matching, as pointed out in Kato (2025c), which extends the argument in Lin et al. (2023).
@@ 3.4 UKL-Riesz Regression
Next, we consider a KL-divergence-motivated convex function. Let $C<\inf_{x}|\alpha(x)|$ be a constant. We define
@@
Under this choice of $g$, the Bregman divergence objective is given as follows [^2] This Bregman divergence objective is derived as follows: $\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)$ $={\mathbb{E}}\Big[-\left(|\alpha(X)|-C\right))\log\left(|\alpha(X)|-C\right)+|\alpha(X)|+\operatorname{sign}\left(\alpha(X)\right)\alpha(X)\log\left(|\alpha(X)|-C\right)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\Big(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\Big].$ :
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)\coloneqq\mathbb{E}\Big[C\log\left(|\alpha(X)|-C\right)+|\alpha(X)|-m\Big(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\Big].$
@@
$\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)$ $=\frac{1}{n}\sum^{n}_{i=1}\left(C\log\left(|\alpha(X_{i})|-C\right)+|\alpha(X_{i})|-m\Big(W_{i},\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\right).$
@@ 3.5 BKL-Riesz Regression
We introduce the BKL divergence and BKL-Riesz, which are motivated by the KL divergence and logistic regression. Let $C<\inf_{x}|\alpha(x)|$ be a constant. We define
@@ 3.6 BP-Riesz Regression
Basu’s power (BP) divergence bridges the squared loss and KL divergence (Basu et al., 1998). Let $C<\inf_{x}|\alpha(x)|$ be a constant. Based on the BP divergence, we introduce the following function:
$g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-C\big)^{1+\delta}-\big(|\alpha|-C\big)}{\delta}-|\alpha|.$
@@
$\partial g^{\text{BP}}(\alpha)=\left(1+\frac{1}{\delta}\right)\operatorname{sign}(\alpha)\Big(\big(|\alpha|-C\big)^{\delta}-1\Big).$
@@
Under this choice of $g$, the Bregman divergence objective is given as follows [^4] This Bregman divergence objective is derived from $\text{BD}_{g^{\text{BP}}}\big(\alpha\big)$ $\coloneqq{\mathbb{E}}\Bigg[-\frac{\big(|\alpha(X)|-C\big)^{1+\delta}-\big(|\alpha(X)|-C\big)}{\delta}+|\alpha(X)|+(1+1/\delta)|\alpha(X)|\Big(\big(|\alpha(X)|-C\big)^{\delta}-1\Big)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,(1+1/\delta)\operatorname{sign}(\alpha)\Big(\big(|\alpha(\cdot)|-C\big)^{\delta}-1\Big)\right)\Bigg].$ :
$\text{BD}_{g^{\text{BP}}}\big(\alpha\big)\coloneqq{\mathbb{E}}\Bigg[\frac{C\big(\big(|\alpha(X)|-C\big)^{\delta}-1\big)}{\delta}+C|\alpha(X)|\big(|\alpha(X)|-C\big)^{\delta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,\left(1+\frac{1}{\delta}\right)\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-C\big)^{\delta}-1\Big)\right)\Bigg].$
@@
$\widehat{\text{BD}}_{g^{\text{BP}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Bigg(\frac{C\big(\big(|\alpha(X_{i})|-C\big)^{\delta}-1\big)}{\delta}+C\big(|\alpha(X_{i})|-C\big)^{\delta}+\big(|\alpha(X_{i})|-C\big)^{1+\delta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W_{i},\left(1+\frac{1}{\delta}\right)\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-C\big)^{\delta}-1\Big)\right)\Bigg).$
Basu’s power divergence bridges the squared loss and the (U)KL divergence. When $\delta=1$, BP-Riesz regression reduces to SQ-Riesz regression, while when $\delta\to 0$, BP-Riesz regression reduces to UKL-Riesz regression. This follows because
$\lim_{\delta\to 0}\frac{\big(|\alpha|-C\big)^{\delta}-1}{\delta}=\log\big(|\alpha|-C\big).$
@@ 3.7 PU-Riesz Regression
We introduce PU learning loss and PU-Riesz, which are motivated by PU learning. Let $C<\inf_{x}|\alpha(x)|$ be some constant. We define $g^{\text{PU}}$ as
@@
PU learning is a classical problem. For example, Lancaster & Imbens (1996) studies this problem under a stratified sampling scheme (Wooldridge, 2001). du Plessis et al. (2015) rediscovers this formulation and calls it unbiased PU learning. Kato et al. (2019) points out the relationship between PU learning and density ratio estimation, and Kato & Teshima (2021) shows that PU learning is a special case of density ratio model fitting under a Bregman divergence. Our results further generalize these results. Note that PU learning in these settings and our setting is called case-control PU learning. There is also another formulation called censoring PU learning (Elkan & Noto, 2008). Kato et al. (2025) considers ATE estimation in a PU learning setup and applies our method in their study.
@@ 4 Automatic Covariate Balancing
Under specific choices of the Riesz representer model and the Bregman divergence, we can guarantee an automatic covariate balancing property. The key tool is the duality between Bregman divergence minimization and covariate balancing methods. This automatic covariate balancing property yields the automatic Neyman orthogonalization property discussed in Section 5, which in turn automatically guarantees Neyman orthogonality for IPW-type estimators. We note that the covariate balancing property does not hold if we use cross fitting.
@@ 4.1 Generalized Linear Models
We can improve estimation accuracy by incorporating additional modeling assumptions. For example, in ATE estimation, we can approximate the propensity score $e_{0}(Z)=P(D=1\mid Z)$ by a logistic model,
where ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function, and $\bm{\beta}$} is the corresponding parameter. Note that in this case, we consider a basis function that receives $Z$ not $X$, or we can interpret that ${\bm{\phi}}(D,Z)$ only depends on $Z$ and is independent of $D$. Plugging this propensity score model into the Riesz representer for ATE, we can approximate the Riesz representer by
$\alpha^{\text{ATE}}_{{\bm{\beta}}}(X)\coloneqq\frac{D}{e_{{\bm{\beta}}}(Z)}+\frac{1-D}{1-e_{{\bm{\beta}}}(Z)}.$
@@
$\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\left(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\right)$
$=\frac{D}{e_{{\bm{\beta}}}(Z)}-\frac{1-D}{1-e_{{\bm{\beta}}}(Z)}$
$=D\Big(1+\exp\Big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big)-(1-D)\Big(1+\exp\Big({\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big).$ (3)
Note that we can also model the Riesz representer as
$\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\left(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\right)$
$=\frac{D}{e_{{\bm{\beta}}}(Z)}-\frac{1-D}{1-e_{{\bm{\beta}}}(Z)}$
$=D\Big(1+\exp\Big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big)-(1-D)\Big(1+\exp\Big({\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big),$ (4)
with including $D$ in the basis function. The choice of basis functions depend on the heterogeneity of $\gamma_{0}(X)$. As we discuss in Section 5, if $\gamma_{0}(x)$ is constant for all $x$, (4.1) may be more appropriate. In contrast, if $\gamma_{0}(x)$ varies across $x$, (4.1) may be more appropriate.
@@ 4.2 Automatic Covariate Balancing
This section presents a covariate balancing property for the basis function ${\bm{\phi}}(X)$ when the Riesz representer model has the form $\alpha(X)=\zeta^{-1}\big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)$. Related properties have been reported in Ben-Michael et al. (2021). More specifically, for Riesz regression, Bruns-Smith & Feller (2022) and Bruns-Smith et al. (2025) discuss the duality between Riesz regression and stable balancing weights (Zubizarreta, 2015). Zhao (2019) also reports the duality between tailored loss minimization and entropy balancing weights. Tarr & Imai (2025) investigates such a duality in covariate balancing method using the support vector machine (SVM). Thus, the duality has garnered attention in covariate balancing (Tan, 2019; Orihara et al., 2025). In density ratio estimation, such a duality is also well known, as discussed in Sugiyama et al. (2007) and Sugiyama et al. (2011b). In the density ratio estimation literature, the resulting balancing property is typically interpreted as moment matching. This study generalizes these results and provides new findings on the balancing property. For the proof, see Appendix B.
@@
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right);$
that is, $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)$ is linear in $\widetilde{g}\left(X_{i},\phi_{1}(X_{i})\right),\dots,\widetilde{g}\left(X_{i},\phi_{p}(X_{i})\right)$. Consider a Riesz representer estimator $\widehat{\alpha}=\alpha_{\widehat{{\bm{\beta}}}}$ trained by generalized Riesz regression (empirical risk minimization) with $\ell_{a}$-penalty as
$\widehat{{\bm{\beta}}}\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha_{{\bm{\beta}}}(X_{i}))+\alpha_{{\bm{\beta}}}(X_{i})\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)-m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}\right\}.$
In this case $\widehat{\alpha}$ satisfies
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\widetilde{g}\big(X_{i},\phi_{j}(X_{i})\big)-m(W_{i},\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right))\Big)\right|\leq\lambda\left|\widehat{\beta}_{j}\right|^{a-1}.$
@@
In addition, if $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$ holds, the following result holds.
Corollary 4.2.
Assume the conditions in Theorem 4.1. Then, if $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$ holds, then the inequality can be written as
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)\right|\leq\lambda\left|\widehat{\beta}_{j}\right|^{a-1}.$
These results correspond to covariate balancing in ATE estimation and moment matching in density ratio estimation.
Examples.
For simplicity, we consider $\ell_{1}$-penalty in generalized Riesz regression. Then, in ATE estimation, if we use a linear link and SQ-Riesz regression, we have
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(D_{i},Z_{i})-\frac{1}{n}\sum^{n}_{i=1}\Big(\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\Big)\right|<\lambda\quad(\text{for}\ j=1,2,\dots,p).$
@@
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\widehat{\alpha}(X_{i})\phi_{j}(Z_{i})-\mathbbm{1}[D_{i}=0]\widehat{\alpha}(X_{i})\phi_{j}(Z_{i})\Big)\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
@@
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-\frac{1}{n}\sum^{n}_{i=1}\phi_{j}(X_{i})\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
@@
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-1\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$